Solve: 3[2x + 2]-7=5
Answer by oscargut(2103)

Solve: 3[2x + 2]-7=5
Solution:
3[2x + 2]-7=5
6x+6-7 = 5
6x = 5-6+7
6x = 6
x = 1
Answer: x = 1

If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com